# Let’s Verify Math Questions Step by Step

### v1

@@ Abstract
Large Language Models (LLMs) have recently achieved remarkable progress in mathematical reasoning. To enable such capabilities, many existing works distill strong reasoning models into long chains of thought or design algorithms to construct high-quality math QA data for training. However, these efforts primarily focus on generating correct reasoning paths and answers, while largely overlooking the validity of the questions themselves. In this work, we propose Math Question Verification (MathQ-Verify), a novel five-stage pipeline designed to rigorously filter ill-posed or under-specified math problems. MathQ-Verify first performs format-level validation to remove redundant instructions and ensure that each question is syntactically well-formed. It then formalizes each question, decomposes it into atomic conditions, and verifies them against mathematical definitions. Next, it detects logical contradictions among these conditions, followed by a goal-oriented completeness check to ensure the question provides sufficient information for solving. To evaluate this task, we use existing benchmarks along with an additional dataset we construct, containing 2,147 math questions with diverse error types, each manually double-validated. Experiments show that MathQ-Verify achieves state-of-the-art performance across multiple benchmarks, improving the F1 score by up to 25 percentage points over the direct verification baseline. It further attains approximately 90% precision and 63% recall through a lightweight model voting scheme. MathQ-Verify offers a scalable and accurate solution for curating reliable mathematical datasets, reducing label noise and avoiding unnecessary computation on invalid questions. Our code and data are available at https://github.com/scuuy/MathQ-Verify.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated impressive performance across a wide range of tasks [16, 10, 26]. In particular, when equipped with chain-of-thought (CoT) of varying lengths [36], some models exhibit the ability to solve moderately complex mathematical problems, such as those from the AIME competition [24]. This reasoning capability is largely attributed to high-quality data sources and efficient training frameworks [9].
@@
However, most existing large-scale mathematical question-answering datasets are composed primarily of synthetically generated QA pairs. A critical issue in such datasets is that if the question itself is flawed, such as being ill-posed or logically inconsistent, then the answer cannot be correct—underscoring the fundamental importance of question correctness.
@@
While some recent efforts have focused on verifying the correctness of answers [32, 41, 33], they often neglect to validate the questions themselves, implicitly assuming that all provided problems are well-formed and mathematically sound. This assumption frequently breaks down in real-world or synthetic data scenarios, where generated problems may include internal contradictions or violate foundational mathematical principles [42]. In such cases, the correctness of the question is uncertain, posing a significant challenge for downstream reasoning tasks. They face the following challenges:
(1) Lack of comprehensive question validation methods.
@@
While several recent studies have begun to examine the validity of mathematical questions [19, 17], their focus is typically limited to a narrow set of error types—such as missing assumptions or vague premises. These efforts fall short of establishing a systematic and comprehensive framework for identifying ill-posed or flawed problems. As a result, many QA datasets, particularly those generated synthetically, still include questions that suffer from internal inconsistencies, logical contradictions, or violations of fundamental mathematical principles.
(2) Lack of stepwise and high-difficulty benchmarks for question verification.
@@
Existing benchmarks, such as MathClean [19], do not provide sufficiently challenging questions, nor do they include fine-grained, stepwise annotations needed to evaluate each stage of a multi-step question verification pipeline. This limits the ability to rigorously assess models’ capabilities in detecting and reasoning about complex flaws in mathematical problem formulations.
@@
To systematically address the challenge of question validity, we propose a structured evaluation framework named MathQ-Verify, which comprises five sequential verification stages: (1) Contaminated Instruction Detection, (2) Linguistic Error Detection, (3) Atomic Condition Error Detection, (4) Cross-condition Conflict Detection, and (5) Condition Completeness Validation. These stages collectively ensure a robust and comprehensive assessment of question quality. An ablation study further confirms the necessity and effectiveness of each individual component within the pipeline.
To enable rigorous evaluation of our pipeline, we introduce a stepwise annotated benchmark dataset, ValiMath, which extends MathClean by incorporating more diverse and challenging examples. ValiMath is constructed from filtered synthetic data derived from the NuminaMath dataset [18], consisting of 2,147 math problems sampled from an initial pool of 10,000 automatically generated questions. Each sample is annotated with fine-grained, stepwise validity labels, encompassing both well-formed and flawed cases to support comprehensive model evaluation across all verification stages.
To summarize, our main contributions are summarized as follows:
  • We construct a new dataset ValiMath, by integrating synthetic problems from NuminaMath and enriching them with structured stepwise labels, specifically designed to support comprehensive evaluation of the correctness of mathematical questions.
  • We propose MathQ-Verify, a pipeline that progressively verifies the correctness of math questions by decomposing them into structured components and checking each part against formalized criteria. MathQ-Verify achieves state-of-the-art results on the two evaluation sets of MathClean, and improves F1 by nearly 15% on our ValiMath compared to direct verification baseline.
@@
2 Related Work
LLMs and Mathematical Reasoning.
@@ 2 Related Work
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of NLP tasks, often surpassing human expert performance in knowledge-intensive domains [7, 21], and generating creative content in writing and design [12, 22]. Beyond general-purpose tasks, recent efforts have increasingly focused on enhancing the reasoning abilities of LLMs, aiming to equip models with deeper, more human-like thinking processes for complex problem solving [28, 39, 31]. Mathematics, as a domain that demands rigorous logical reasoning, has become a popular benchmark for evaluating these capabilities. Commercial products like ChatGPT [25], Claude [4] and Gemini [34] have shown strong performance on math evaluations, while open-source models such as DeepSeek-R1 [9] and the Qwen3 series [5] have also made competitive advances, highlighting the importance of training objectives like reinforcement learning. These advancements have been driven by the availability of high-quality reasoning datasets, especially in math problem solving [20, 8, 2, 14]. Our work enhances this trend by proposing a novel approach to filter and verify math problems, improving data reliability for training and evaluation.
Synthetic Data.
@@
Training LLMs requires massive amounts of data [38, 1], but existing web corpora have become increasingly insufficient to support full training pipelines, particularly in the pretraining stage [35, 29]. In addition, many open-source datasets suffer from noisy supervision and inconsistent quality, making them suboptimal for model training [15]. To address this, recent work has explored synthetic data generation using LLMs [23, 30], enabling controllable data distributions and customizable supervision signals. However, synthetic data especially in the domain of mathematical question generation often contains subtle or critical errors in both questions and solutions, highlighting the need for rigorous verification before such data can be reliably used for training or evaluation. Based on this, we explore methods for verifying and filtering erroneous mathematical data.
Verification in Math QA.
@@
Much attention has been given to improving the accuracy of LLMs on math problem answering, including developing better reasoning strategies and addressing common model failure modes [37, 40]. Models fine-tuned with reinforcement learning have even demonstrated the ability to solve competition-level math problems [9, 25, 6]. However, when presented with flawed or ill-posed questions, LLMs often lack the ability to detect inconsistencies or exhibit overthinking behaviors [27, 11]. Recent efforts such as MathClean [19] and QuestBench [17] have introduced benchmarks for identifying flawed problems and missing assumptions, revealing substantial gaps in model robustness. Despite these advances, structured and comprehensive verification for math problems has yet to receive broad attention. To bridge this gap, our work proposes a progressive, formalized verification pipeline that identifies and filters flawed problems based on a structured decomposition of the problem semantics.
3 Methodology
@@
In this section, we introduce our MathQ-Verify, as shown in Figure 1, designed to verify the correctness of synthetic questions. Our methodology systematically deconstructs each question into its fundamental components and evaluates them step by step through a carefully designed pipeline. This pipeline includes multiple stages, including structural analysis, logical consistency checks, and contextual validation, to ensure a comprehensive assessment of the question’s quality and reasonableness. By rigorously analyzing and verifying each question, our approach aims to establish a robust framework for evaluating the validity and soundness of mathematical problems.
3.1 Problem Formulation
@@
We define the set of questions in the mathematical QA dataset as $\mathcal{Q}=\{q_{i}\}_{i=1}^{L}$, where each $q_{i}$ denotes a natural language math question. This set can be partitioned into two disjoint subsets: a set of well-formed (valid) questions $\mathcal{Q}^{\text{valid}}$ and a set of ill-formed (invalid) questions $\mathcal{Q}^{\text{invalid}}$, such that:
$Q=Q_{\text{valid}}\cup Q_{\text{invalid}},\quad Q_{\text{valid}}\cap Q_{\text{invalid}}=\emptyset$
@@
Identifying whether a given question is well-formed is crucial for downstream tasks such as automated solving and dataset refinement. Therefore, we aim to learn a verification function $V:\mathcal{Q}\rightarrow\{0,1\}$, where $V(q_{i})=1$ indicates that $q_{i}$ is a valid question, and $V(q_{i})=0$ otherwise.
@@
To determine whether a question $q_{i}$ is valid, we propose to decompose it into two structured components, which are then used as the basis for verification through $V(q_{i})$. Specifically:
@@
Atomic Conditions, $\mathcal{P}(q_{i})=\{P_{1},P_{2},\dots,P_{N}\}$: These represent the foundational mathematical statements embedded within the question, serving as the building blocks for its logical structure.
@@
Figure 1: An overview of our MathQ-Verify framework. Given a math question, we extract atomic conditions $P_{i}$ and the goal $G_{i}$, and conduct a five-step verification process: (1) detection of contaminated instructions, (2) linguistic error detection, (3) atomic condition verification, (4) cross-condition conflict detection, and (5) condition completeness validation. Only questions that pass all checks are retained as correct structured math questions.
3.2 Question Correctness Verification
As outlined in the previous section, verifying the correctness of a mathematical question, denoted as $V(q_{i})$, requires a systematic approach to deconstruct and analyze its components. In this section, we introduce the five stages of the Math Question Verification (MathQ-Verify) pipeline, which is specifically designed for this purpose. Let’s now verify the math question step by step.
Step1: Contaminated Instruction Detection.
To construct a high-quality textual math QA dataset, the first step is to detect and remove contaminated instructions, such as those containing misleading instructions or leaked answers. This process ensures that only semantically and syntactically appropriate items are retained for downstream tasks. To this end, we define a binary instruction validity indicator for each question $q_{i}$ as:
$\text{InstValid}(q_{i})=1\iff\text{IsMathQ}(q_{i})\land\neg\text{HasMisleadingCue}(q_{i})\land\neg\text{HasAnswerLeak}(q_{i})$
This holds if and only if $q_{i}$ is a genuine math question, free from misleading linguistic patterns and explicit answer leakage. The indicator reflects whether the instruction adheres to the structural and semantic criteria expected of high-quality mathematical queries. In particular, we perform a model-centric inspection that assesses both the mathematical nature and instructional integrity of each input. For example, if a question $q_{i}$ includes textual cues such as “please rewrite” or “please rephrase”, or if it appends an explicit answer within the same prompt, it is considered misaligned with the problem-solving objective. Such artifacts may distract the model from focusing on solving the mathematical content and are thus filtered out during this validation phase.
Step2: Linguistic Error Detection.
After identifying and isolating the core mathematical question in the first step, we proceed to detect linguistic imperfections that may hinder readability or model interpretation. This step focuses on linguistic-level issues such as spelling mistakes, grammatical errors, and LaTeX formatting anomalies. While these can be addressed with rule-based heuristics, we adopt the Qwen-2.5-7B-Instruct model for simplicity and generality.
Following this, we obtain a set of clean math problems, where "clean" refers to questions that are exclusively mathematical in form, regardless of clarity or correctness. Formally, a question $q_{i}$ is clean if it contains no non-semantic errors that affect surface quality, and its content can be decomposed into a set of mathematical atomic conditions $\mathcal{P}(q_{i})$ and a target query $\mathrm{Query}(q_{i})$.
$\text{Clean}(q_{i})=1\iff\neg\,\text{LinguisticError}(q_{i})\land\exists\,\mathcal{P}(q_{i}),\,\mathrm{Query}(q_{i})$
Step3: Atomic Condition Error Detection.
For a clean mathematical question $q_{i}$, we systematically examine each of its atomic conditions $\{P_{1},P_{2},\ldots,P_{N}\}$ to ensure their consistency with established mathematical principles. A atomic condition is defined as a foundational statement that encapsulates a distinct mathematical fact (e.g., $x\in\mathbb{Z}$). Each condition $P_{j}$ undergoes rigorous validation to confirm its adherence to the rules of the corresponding mathematical domain. To guarantee validity, we verify the following for all conditions $P_{j}\in P(q_{i})$:
$\text{Valid}(q_{i})=1\iff\forall P_{j}\in\mathcal{P}(q_{i}),\quad\text{Valid}(P_{j})=1$
While some flexibility may be allowed for commonly understood concepts or heuristic steps (often referred to as "magic steps") that aid in mathematical reasoning, any condition that contradicts fundamental definitions is strictly rejected. For example, defining a discrete quantity as continuous would be considered invalid. To illustrate, consider the following question, which includes a atomic condition that violates core mathematical principles:

### v2

@@ Abstract.
Large Language Models (LLMs) have recently achieved remarkable progress in mathematical reasoning. To enable such capabilities, many existing works distill strong reasoning models into long chains of thought or design algorithms to construct high-quality math question-answer (QA) data for training. However, these efforts primarily focus on generating correct reasoning paths and answers, while largely overlooking the correctness of the questions themselves. In this work, we present ValiMath, a benchmark consisting of 2147 human-verified mathematical questions covering a wide range of domains such as arithmetic, algebra, and geometry, which are synthesized and curated from the NuminaMath dataset. Each question is annotated with its logical structure, domain coverage, and question correctness, enabling fine-grained evaluation of question quality. ValiMath serves as a high-quality gold-standard test set for validating mathematical questions in LLM training corpora. Building upon this benchmark, we further propose MathQ-Verify, a pipeline that performs fine-grained parsing of mathematical questions into atomic assumptions and conclusions, and evaluates their semantic soundness through consistency checks. This pipeline achieves high precision in detecting flawed questions and provides a reliable foundation for cleaning noisy mathematical datasets. Experiments show that MathQ-Verify achieves state-of-the-art performance across multiple benchmarks, improving the F1 score by up to 25 percentage points over the direct verification baseline. MathQ-Verify offers a scalable and accurate solution for curating reliable mathematical datasets, reducing label noise and avoiding unnecessary computation on invalid questions. Our code and data are available at the repository https://github.com/OpenDCAI/MathQ-Verify.
Data Cleaning, Math Reasoning, Question Verification
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3785700isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Machine learningccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
Large Language Models have demonstrated impressive performance across a wide range of tasks (Kaddour et al., 2023; Fan et al., 2023; OpenAI et al., 2024b). In particular, when equipped with chain-of-thought (CoT) of varying lengths (Wei et al., 2023), some models exhibit the ability to solve moderately complex mathematical problems, such as those from the AIME competition (of America, 2025). This reasoning capability is largely attributed to high-quality data sources and efficient training frameworks (DeepSeek-AI et al., 2025).
@@
However, most existing large scale mathematical question and answer datasets are composed primarily of synthetically generated QA pairs. A critical issue in such datasets is that if the question itself is flawed, such as being ill-posed or logically inconsistent, then the answer cannot be correct—underscoring the fundamental importance of question correctness.
@@
While some recent efforts have focused on verifying the correctness of answers (Song et al., 2025; Zheng et al., 2024; Sun et al., 2025), they often neglect to validate the questions themselves, implicitly assuming that all provided problems are well-formed and mathematically sound. This assumption frequently breaks down in real-world or synthetic data scenarios, where generated problems may include internal contradictions or violate foundational mathematical principles (Zhou et al., 2024). In such cases, the correctness of the question is uncertain, posing a significant challenge for downstream reasoning tasks. They face the following challenges:
@@
(1) Lack of comprehensive question validation methods. While several recent studies have begun to examine the correctness of mathematical questions (Liang et al., 2025; Li et al., 2025), their focus is typically limited to a narrow set of error types—such as missing assumptions or vague premises. These efforts fall short of establishing a systematic and comprehensive framework for identifying ill-posed or flawed problems. As a result, many QA datasets, particularly those generated synthetically, still include questions that suffer from internal inconsistencies, logical contradictions, or violations of fundamental mathematical principles.
@@
(2) Lack of stepwise and high-difficulty benchmarks for question verification. Existing benchmarks, such as MathClean (Liang et al., 2025), do not provide sufficiently challenging questions, nor do they include fine-grained, stepwise annotations needed to evaluate each stage of a multi-step question verification pipeline. This limits the ability to rigorously assess models’ capabilities in detecting and reasoning about complex flaws in mathematical problem formulations.
To advance the rigorous evaluation of mathematical question correctness, we present ValiMath, a stepwise annotated benchmark dataset specifically designed to support fine-grained analysis of question formulation quality. ValiMath extends the existing MathClean benchmark by introducing a broader range of synthetically generated yet realistically flawed math problems, filtered and curated from the NuminaMath dataset (Li et al., 2024). From an initial pool of 10,000 auto-generated questions, we sample 2,147 representative problems and annotate each with a sequence of structured correctness labels. These labels span multiple levels of semantic and logical fidelity, capturing both well-formed and defective cases. By offering explicit annotations at each verification step, ValiMath enables systematic and interpretable evaluation of question generation pipelines, and provides a foundation for benchmarking future efforts in automatic data validation.
@@
Building on ValiMath, we further propose MathQ-Verify, a structured verification framework that operationalizes the evaluation of question correctness through five sequential stages: (1) Contaminated Instruction Detection, (2) Linguistic Error Detection, (3) Atomic Condition Error Detection, (4) Cross-condition Conflict Detection, and (5) Condition Completeness Validation. Each stage targets a specific class of flaws commonly observed in synthetic question datasets. Through comprehensive ablation studies, we demonstrate the necessity and effectiveness of each component, establishing MathQ-Verify as a robust reference pipeline for quality assessment in math question generation tasks. In summary, our contribution are as follows:
  • We construct a new benchmark dataset, ValiMath, which supports fine-grained evaluation of mathematical question correctness. ValiMath contains 2,147 math problems curated from synthetic sources and annotated with structured, stepwise correctness labels that cover a wide spectrum of realistic formulation errors. It provides a principled foundation for evaluating data validation systems in the mathematical reasoning domain.
  • We design a new verification pipeline, MathQ-Verify, which decomposes the validation process into five interpretable stages aligned with formalized criteria. This modular pipeline significantly improves validation performance, achieving state-of-the-art results on MathClean and yielding a 15% absolute F1 gain over direct baselines on ValiMath.
@@
2. Related Work
@@
LLMs and Math Reasoning. Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of NLP tasks, often surpassing human expert performance in knowledge-intensive domains (Brodeur et al., 2024; Luo et al., 2024), and generating creative content in writing and design (Franceschelli and Musolesi, 2024; Ma et al., 2024). Beyond general-purpose tasks, recent efforts have increasingly focused on enhancing the reasoning abilities of LLMs, aiming to equip models with deeper, more human-like thinking processes for complex problem solving (Patil, 2025; Yang et al., 2025; Ren et al., 2025). Mathematics, as a domain that demands rigorous logical reasoning, has become a popular benchmark for evaluating these capabilities (Saxton et al., 2019; Lewkowycz et al., 2022; Drori et al., 2022). Commercial products like ChatGPT (OpenAI et al., 2024a), Claude (Anthropic, 2024) and Gemini (Team et al., 2024) have shown strong performance on math evaluations, while open-source models such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and the Qwen3 series (Bai et al., 2023) have also made competitive advances, highlighting the importance of training objectives like reinforcement learning (Bansal et al., 2019; Fawzi et al., 2022; Zhao et al., 2023). These advancements have been driven by the availability of high-quality reasoning datasets, especially in math problem solving (Lu et al., 2024a; Cobbe et al., 2021; Albalak et al., 2025; Hendrycks et al., 2021). Our work enhances this trend by proposing a new benchmark to evaluate the models’ capability of filtering and verifying the potential incorrect math questions.
@@
Synthetic Data. Training LLMs requires massive amounts of data (Yang et al., 2024; Abdin et al., 2024), but existing web corpora have become increasingly insufficient to support full training pipelines, particularly in the pretraining stage (Villalobos et al., 2024; perełkiewicz2024reviewchallengesmassivewebmined). In addition, many open-source datasets suffer from noisy supervision and inconsistent quality, making them suboptimal for model training (Jiang et al., 2024). To address this, recent work has explored synthetic data generation using LLMs (Nadas et al., 2025; Qin et al., 2025), enabling controllable data distributions and customizable supervision signals. However, synthetic data especially in the domain of mathematical question generation often contains subtle or critical errors in both questions and solutions (Lu et al., 2024b; Vendrow et al., 2025; Shen et al., 2025; Wang et al., 2023), highlighting the need for rigorous verification before such data can be reliably used for training or evaluation. Based on this, we explore methods as a reference for verifying and filtering erroneous mathematical data.
@@
Verification in Math QA. Much attention has been given to improving the accuracy of LLMs on math problem answering, including developing better reasoning strategies and addressing common model failure modes (Xu et al., 2024; Zhang et al., 2024). Models fine-tuned with reinforcement learning have even demonstrated the ability to solve competition-level math problems (DeepSeek-AI et al., 2025; OpenAI et al., 2024a; Bercovich et al., 2025). However, when presented with flawed or ill-posed questions, LLMs often lack the ability to detect inconsistencies or exhibit overthinking behaviors (Ouyang, 2025; Fan et al., 2025). Recent efforts such as MathClean (Liang et al., 2025) and QuestBench (Li et al., 2025) have introduced benchmarks for identifying flawed problems and missing assumptions, revealing substantial gaps in model robustness. Despite these advances, structured and comprehensive verification for math problems has yet to receive broad attention. To bridge this gap, we introduce ValiMath, a benchmark dataset tailored for stepwise verification of question correctness. To showcase its utility, we provide a formalized multi-stage verification pipeline as a reference approach, illustrating how semantic decomposition enables fine-grained quality assessment.
Figure 1. An overview of our MathQ-Verify framework. Given a math question, we extract atomic conditions $P_{i}$ and the goal $G_{i}$, and conduct a five-step verification process: (1) detection of contaminated instructions, (2) linguistic error detection, (3) atomic condition verification, (4) cross-condition conflict detection, and (5) condition completeness validation. Only questions that pass all checks are retained as correct structured math questions.
3. ValiMath: The Question Verification Bench
3.1. Dataset Overview
While MathClean(Liang et al., 2025) provides a benchmark for validating math question correctness, it primarily contains simpler questions and lacks fine-grained annotations. To introduce a more diverse range of difficulties for comprehensive evaluation, we present ValiMath, a benchmark of 2,147 questions (1,299 correct, 848 incorrect) covering five diverse error types (defined in Section 4) as shown in Table 1.
Table 1. Data distribution of ValiMath.
Category Count %
Instruction Error 143 6.66
Nonsemantic Error 119 5.54
Minimal Domain Error 171 7.96
Contradiction Error 300 13.97
Completeness Error 115 5.36
Correct Instances 1299 60.50
Our dataset encompasses a wide range of difficulty levels and diverse mathematical domain labels. The correctness of the data is ensured through a rigorous filtering and validation process. We have reorganized and developed a new secondary classification system based on the MSC classification standard (American Mathematical Society and zbMATH Open, 2020), which has been systematically applied to our data. The dataset is constructed through a meticulous multi-stage pipeline, involving mathematics domain experts to guarantee high-quality annotations. This comprehensive process ensures both the correctness of the questions and the accuracy of the labels. More details and the complete data construction process are provided as follows.
3.2. The Construction of ValiMath
This subsection introduces the construction process of the ValiMath.
Step1: Data Synthesis. For synthetic data in the mathematics domain, especially for the generation of questions, there is a high demand for diversity. Additionally, the synthetic data itself tends to align more closely with the distribution of model training data. To construct a diverse set of questions, while also introducing more challenging problems compared to MathClean (Liang et al., 2025), we thoroughly researched existing prompts used for generating synthetic questions. Based on our findings, we carefully designed a prompt (provided in repository) that ensures the diversity of the generated questions. To ensure the generated mathematical problems possess a certain level of difficulty and challenge, we ultimately used NuminaMath (Li et al., 2024) as our source data for question synthesis.
Furthermore, to capture a broad range of potential errors that the model might generate, we employed models of different sizes for question synthesis. Stronger reasoning models tend to generate problems with a lower error rate during normal question synthesis, while smaller models are more likely to produce erroneous questions (Wei et al., 2022; Honovich et al., 2022; Lu et al., 2024b; Chowdhery et al., 2022). Therefore, we used a variety of models, including: Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, GPT-4o, and GPT-o3mini, ensuring that each model generated a similar number of questions. In total, we synthesized 10,000 mathematical questions in this stage.
Step2: Filtering. To ensure the uniqueness and determinacy of the final answers, we applied a GPT-4o-based prompt filtering process to exclude all multiple-choice questions. It is important to note that, since our objective is to construct a dataset containing both correct and incorrect questions, no additional filtering was performed beyond this step. This preserves the natural distribution of model-generated outputs.
Step3: Data Selection. To reduce annotation costs in advance, we observed during preliminary experiments that general-purpose models such as GPT-4o tend to generate more correct than incorrect mathematical questions in our raw data (with fewer than 30% incorrect samples). To address this, we first performed a coarse-grained filtering step to retain a higher proportion of incorrect questions for annotation.
Specifically, we employed Qwen2.5-72B-Instruct to directly assess the correctness of each generated question using the baseline judgment method described in the main paper. Based on these results, we curated a subset of approximately 4,000 mathematical questions with a target ratio of correct to incorrect samples set at 1:3. This step effectively reduced downstream annotation workload while preserving as many erroneous questions as possible, thereby enhancing the effectiveness of subsequent data verification.
Step4: Expert Annotation. After obtaining the initial set of 4,000 mathematical questions, we invited domain experts in mathematics to perform detailed annotations. These annotators were selected through a three-stage screening process from an initial pool of approximately 150 candidates. They were evaluated based on the correctness of their annotations on 10, 50, and 100 questions, respectively. Ultimately, 8 experts were chosen. After a month of intensive training, these experts fully met our standards for annotation quality and domain expertise. In addition, their ethical integrity and moral standards were also assessed and confirmed to meet our predefined requirements. The goal was to ensure the accuracy of the binary correctness labels (”correct” or ”incorrect”) and, for incorrect questions, to further annotate the specific steps where errors occurred, following the five-stage framework defined in MathQ-Verify. The annotation process proceeded as follows:
  1. (1)
    Annotation Guidelines. Experts were provided with comprehensive annotation instructions (detailed in Appendix B), including precise definitions of each step in the MathQ-Verify process as outlined in the main paper. Each question was evaluated across the five reasoning stages, and labeled accordingly as either “correct” or “incorrect.” For incorrect questions, annotators further specified the earliest step at which the error occurred.
  2. (2)
    Independent Dual Annotation. Each question was independently annotated by at least two domain experts to mitigate individual biases. Discrepancies between annotations were resolved through a final joint review process, during which experts reached a consensus based on predefined labeling criteria.
  3. (3)
  4. (4)
    Subjectivity Analysis. To further assess and control for annotator subjectivity, we randomly sampled a subset of annotated questions for inter-annotator agreement analysis. We applied standard consistency metrics to quantify labeling stability across experts and identify any systemic annotation biases.
  5. (5)
    Quality Assurance. Through the above multi-step annotation workflow, we ensured the reliability, consistency, and accuracy of the labeled dataset, providing a high-quality foundation for dataset construction and downstream evaluation.
4. MathQ-Verify
@@
In this section, we introduce our MathQ-Verify, as shown in Figure 1, designed to verify the correctness of mathematical questions. Our methodology systematically deconstructs each question into its fundamental components and evaluates them step by step through a carefully designed pipeline. This pipeline includes multiple stages, including structural analysis, logical consistency checks, and contextual validation, to ensure a comprehensive assessment of the question’s quality and reasonableness. By rigorously analyzing and verifying each question, our approach aims to establish a robust framework for evaluating the correctness and soundness of mathematical questions.
4.1. Problem Formulation
@@
We define the set of questions in the mathematical QA dataset as $\mathcal{Q}=\{q_{i}\}_{i=1}^{L}$, where $L$ is the number of questions and each $q_{i}$ denotes a natural language math question. A valid question, called correct question as well, means it has clear conditions, is well-defined and free from contradictions, and can be solved, and the invalid question is on the contrary side. This set can be partitioned into two disjoint subsets: a set of well-formed (valid) questions $\mathcal{Q}^{\text{valid}}$ and a set of ill-formed (invalid) questions $\mathcal{Q}^{\text{invalid}}$, such that:
$\mathcal{Q}=\mathcal{Q}_{\text{valid}}\cup\mathcal{Q}_{\text{invalid}},\quad\mathcal{Q}_{\text{valid}}\cap\mathcal{Q}_{\text{invalid}}=\emptyset$
@@
Identifying whether a given question is valid is crucial for downstream tasks such as dataset refinement. Therefore, we aim to learn a verification function $V:\mathcal{Q}\rightarrow\{0,1\}$, where $V(q_{i})=1$ indicates that $q_{i}$ is a valid question, and $V(q_{i})=0$ otherwise.
@@
To determine whether a question $q_{i}$ is valid, we propose to decompose it into two structured components using LLM (detailed in repository), which are then used as the basis for verification through $V(q_{i})$. Specifically:
@@
Atomic Conditions, $\mathcal{P}(q_{i})=\{P_{1},P_{2},\dots,P_{N}\}$: These represent the minimal and indivisible conditions in the question, serving as the building blocks for a complete math question.